# CITADEL: A Semi-Supervised Active Learning Framework for Malware Detection Under Continuous Distribution Drift

### v1

@@ III-B Motivating Example
A core challenge in malware detection is the emergence of new variants that introduce previously unseen or rare semantic features, leading to concept drift. Traditional supervised learning models are typically trained on static, labeled datasets and often fail to generalize when confronted with novel traits at test time. Compounding this issue, producing high-quality labels for malware is both time-consuming and resource-intensive—prior work shows that vendors can take weeks to reach consensus on newly discovered variants [18, 19]. Moreover, malware analysts can manually review only about 80 samples per day [17], leaving the vast majority of malware unlabeled. This creates an untapped opportunity to utilize the abundance of unlabeled samples in the detection pipeline [21]. Motivated by this observation, we explore SSL as a means to both detect and adapt to concept drift by jointly leveraging labeled and unlabeled data. In particular, we investigate the FixMatch framework [22], which has demonstrated strong performance in other domains by effectively utilizing unlabeled samples.
@@
FixMatch combines pseudo-labeling with consistency regularization through weak and strong data augmentation techniques. For an unlabeled input sample , FixMatch generates a pseudo-label by applying a weak augmentation and passing it through the current model. This pseudo-label is then used as a supervisory signal for the corresponding strongly augmented version . The learning objective enforces consistency between the model’s predictions for and , thereby encouraging the model to produce stable predictions under input perturbations. Weak augmentations (e.g., horizontal flips, minor translations, rotations) preserve the semantic content of the input and facilitate reliable pseudo-labeling. Strong augmentations (e.g., color jitter, brightness alteration, cutout) introduce stochastic feature perturbations that can mask or distort key input features [50]. Despite this, the model is trained to maintain prediction consistency, thereby forcing it to generalize beyond specific feature dependencies.
@@
In the context of malware detection, this mechanism is particularly useful for learning from drift-induced features. Let be a malware sample containing both a known semantic feature and a novel, previously unseen feature . The weak augmentation retains both features, enabling the model to generate a high-confidence pseudo-label based on . If the strong augmentation masks via stochastic feature perturbation, the detector must rely on to maintain consistency with , thereby incorporating into the learned decision boundary. This implicit adaptation supports concept drift handling by progressively integrating novel features without explicit annotation. Prior work demonstrates that this mechanism promotes the learning of low-density decision boundaries and improved feature alignment between labeled and unlabeled samples [50].
@@
In our setting, FixMatch enables the model to bridge concept drift by exploiting the structure of unlabeled data, thereby facilitating more robust detection of unseen or evolving malware variants. The training dataset at step in the Figure 3 is divided into two sets labeled, at step and unlabeled at step . Let denote the labeled dataset and the unlabeled dataset. The overall FixMatch objective is a weighted sum of the supervised loss on labeled data and the unsupervised consistency loss on high-confidence pseudo-labeled data:
where is the standard cross-entropy (CE) loss computed over labeled samples:
@@
Here, denotes the malware detector parameterized by , is the strongly augmented view of the unlabeled input , and is the predicted softmax probability from the weakly augmented version used to obtain pseudo-label . The indicator function filters pseudo-labels based on a confidence threshold . The hyperparameter controls the influence of the unsupervised loss. In our experiments, we set and , following best practices from prior work [22]. This formulation encourages the model to produce consistent predictions across different input perturbations and to generalize from labeled data while gradually adapting to new or shifting semantic features, capabilities essential for effective malware detection under concept drift.
IV Overview of CITADEL
@@
In this section, we introduce our semi-supervised active learning (AL) framework, CITADEL, for adapting to concept drift in malware detection. The overall process is outlined in Algorithm 1 (Semi-supervised Initial Training) and Algorithm 2 (Active Learning).
@@
We first propose two domain-specific augmentations—Bernoulli bit flip and feature mask—to emulate realistic malware evolution and perturb semantic features. As shown in step of Figure 3, the unlabeled dataset is divided into a WeakFlip set and a StrongFlip set . In step , we perform semi-supervised training using a unified objective that combines supervised contrastive loss with the standard semi-supervised loss, enhancing class separation, particularly for ambiguous and drifted samples. To support adaptation under limited labeling budgets, we design a multi-criteria sample selection strategy in step , which guides the AL process at step , based on the budget defined in step . The cycle from steps to is repeated until the new test set becomes available.
IV-A Augmentations for Malware Representations
@@
CITADEL is built upon FixMatch [22], which is originally developed for the computer vision domain, where augmentations such as rotation, translation, and color distortion are semantically meaningful. In contrast, malware classification operates on a semantically rich, high-dimensional tabular feature space. These features are typically derived from static analysis of binaries (e.g., header metadata, byte sequences, and API call usage), and are subject to strong semantic constraints [47]. Consequently, image-based augmentations are not applicable in this domain and often fail to produce valid perturbations. In the case of Android malware, features are commonly represented as static binary vectors, where each dimension indicates the presence or absence of specific behavioral attributes, permissions, or API calls [6]. These binary vectors encode meaningful structural and behavioral properties of malware samples.
To address this challenge, we propose two augmentation strategies (shown in Algorithm 1 line: 6-7) tailored for static binary feature vectors—Bernoulli bit flip and Bernoulli feature mask (see Figure 4). Since there is no deterministic pattern of distribution shifts caused by concept drift in Android malware [51], we adopt a randomized approach to simulate such drift directly in the feature space through these augmentations. These strategies play a role analogous to data transformations in FixMatch by enabling consistency regularization within the feature space. As a result, the model learns feature representations that remain stable under semantically valid perturbations. MORSE [12] also employs FixMatch, but with a different goal, detecting malware in noisy or incorrectly labeled datasets to improve noise tolerance, without addressing concept drift. Their augmentation method introduces noise by replacing feature values across different samples, where a subset of features is selected according to a Bernoulli distribution.
Bernoulli Bit Flip. This augmentation strategy simulates behavioral mutations commonly observed across malware variants by probabilistically flipping individual bits in the binary feature vector. In this representation, a bit value of 1 indicates the presence of a specific API call or code pattern, while 0 indicates its absence. Given an input feature vector , we initially consider injecting noise using a discrete uniform distribution.
where each is sampled independently and uniformly from , and is the input feature dimension. The noise vector is then combined with the original input using bitwise XOR to generate the augmented sample:
Inspired by Wu et al. [12], we adopt Bernoulli distribution to simulate distribution shift using the original input, as it produces independent binary outcome for each feature with a controlled probability , which aligns with the binary representation of malware features.
Here, denotes the bitwise XOR operation, and is a noise vector sampled independently for each bit, where each bit is flipped with probability . Small values of correspond to subtle modifications that emulate minor obfuscations or mutations, while larger values introduce more substantial perturbations, potentially degrading semantic coherence.
@@
Bernoulli Feature Mask. In this augmentation, we intend to simulate code obfuscation by randomly zeroing out a subset of the features, effectively removing observable behaviors from the sample. Let be a binary mask where each bit is zeroed out with probability . The masked input is defined as:
@@
Weak and Strong Augmentation. In CITADEL framework, weak and strong augmentation serve distinct roles. For weak augmentation, we apply a low perturbation probability preserving core semantics of a sample and used for generating confident pseudo-labels, (shown in Algorithm 1 line: 8). For strong augmentation, we apply comparatively high perturbation probability to introduce more noise.
To evaluate the individual contribution of each augmentation strategy, we separately explore the effect of Bernoulli bit flip and Bernoulli feature mask on model generalization and concept drift adaptation. We empirically determine the perturbation rates and to ensure that augmentations are both semantically meaningful and effective for training under concept drift. For Bernoulli bit flip, we apply weak and strong perturbations using flip probabilites and respectively. Similarty, for Bernoulli feature mask, we also apply weak and strong mask with probabilities and , respectively. We also explore a combination of random bit flip and feature masking, and empirically evaluate which perturbation settings provide the best performance under concept drift scenario.
IV-B CITADEL’s Objective Function
@@
The supervised contrastive loss follows the formulation on [25]. Given a batch of encoded feature vectors and their corresponding labels , the loss is defined as:
@@
IV-C Multi-Criteria Sample Selection
Active learning offers a promising approach to reducing labeling costs in malware detection [11], particularly in the presence of concept drift. However, selecting informative samples in this context remains challenging due to the inherent sparsity, class imbalance, and evolving nature of malware behaviors. We empirically observe that no single sampling criterion reliably captures all informative instances under such dynamics. To address this, we introduce a multi-criteria sampling strategy based on three complementary indicators: (1) boundary proximity, quantified using the softmax margin; (2) feature displacement, measured via the -norm distance in embedding space; and (3) prediction uncertainty, captured through low-confidence outputs. Each indicator provides a distinct perspective on sample informativeness. We hypothesize that their combination enables more robust and stable adaptation in the face of continual malware evolution. Algorithm 2 presents the detailed procedure of our active learning method.
Softmax Margin for Boundary Samples. The first sampling criterion leverages the model’s softmax prediction margin, which reflects the proximity of a sample to the decision boundary. For an unlabeled input , let denote the model’s softmax output vector. The softmax margin is defined as the difference between the top two predicted class probabilities:
where and are the highest and second-highest values in , respectively. A smaller margin indicates greater uncertainty and typically corresponds to samples near the model’s decision boundary. Such samples are more likely to be informative when labeled. Margin-based uncertainty sampling is a well-established strategy in the active learning literature [52], and has demonstrated robustness even under distributional shift [53].
-Norm for Feature-Space Drift. While the softmax margin captures local uncertainty near the classifier’s decision boundary, it may fail to identify outliers or samples that are semantically distant from the training distribution—a common scenario under concept drift. To address this, we compute the -norm distance between each unlabeled sample and its nearest labeled neighbor in the model’s learned feature space.
@@ IV-C Multi-Criteria Sample Selection
To select an appropriate value, we empirically evaluated the active learning performance for in increments of 0.1. Our results showed only minor performance variation across values. We ultimately select due to its consistent performance across multiple settings and its theoretical grounding in high-dimensional vector spaces [54, 55].
@@
Samples with low confidence often correspond to ambiguous instances or previously unseen malware classes. Prior work has demonstrated that such samples are strongly correlated with misclassification [56] and can be effectively utilized for active learning under distributional shift [36, 35]. In the context of malware detection, this criterion enables the model to prioritize uncertain or suspicious samples that may not be captured by margin-based or distance-based strategies alone.
Sample Selection Strategy. First, we take the intersection of the three criteria to ensure that we do not select the same samples combining the criteria. Then, we leverage the complementary strengths of the three uncertainty criteria: margin-based (), feature-space distance (), and confidence-based (), we compute a unified selection score for each unlabeled sample. As shown in Algorithm 2, lines 3–6, each score component is first min-max normalized across the unlabeled pool:
Here, denotes the normalized value, and are weighting coefficients (set to 1 by default unless otherwise specified). This formulation assigns higher selection scores to samples with low prediction margins, high feature-space distance, and low model confidence—each of which signals potential informativeness under concept drift. The top- samples with the highest scores are selected for labeling.
The sample selection and retraining procedures are detailed in Algorithm 2, lines 10–13, which invoke the semi-supervised training routine in Algorithm 1. Compared to prior active learning approaches in malware detection that rely on a single uncertainty metric [20, 11], our multi-criteria strategy enables more robust adaptation to drift by jointly targeting decision boundary ambiguity and semantic outliers in the learned feature space.
| Year | APIGraph [13] | Chen-AZ [11] | LAMDA [15] | MaMaDroid [14] | ||||
@@
V Evaluation
In this section, we present the comprehensive evaluation of CITADEL. Section V-A introduces the datasets, followed by Section V-B, which analyzes CITADEL without active learning (AL) under varying labeled ratios and feature augmentations. Section V-C compares CITADEL with baseline methods without AL, while Section V-D extends the comparison to the AL framework. Finally, Section V-E discusses the computational complexity. Additional analysis and experimental results are provided in the Appendix—Appendix A details family and singleton statistics, Appendix B visualizes the performance trends of CITADEL and Chen-AL, Appendix C shows results of hyperparameter tuning, Appendix D examines label noise robustness, Appendix E explores curriculum learning integration, and Appendix F reports results on the EMBER [57].

### v2

@@ 3.2 Motivating Example
A core challenge in malware detection is the continual emergence of new variants that introduce unseen and rare semantic features. Models trained on static, labeled datasets often struggle to generalize to these changes, resulting in degraded performance over time. Compounding this issue, producing high-quality labels for malware is both time-consuming and resource-intensive. Prior studies show that vendors may take weeks to reach consensus on newly discovered variants [zhu2020measuring, kantchelian2015better]. Moreover, malware analysts can manually review only about 80 samples per day [miller2016reviewer], leaving the vast majority of malwares unlabeled.
This imbalance creates creates a large pool of unlabeled samples that is unused by traditional detection pipelines [apruzzese2022sok]. Motivated by this observation, we explore semi-supervised learning (SSL) as a means to improve malware detection under evolving data distributions by jointly leveraging labeled data and abundant unlabeled samples. In particular, we investigate the FixMatch framework [sohn2020fixmatch], which has demonstrated strong performance in settings where labeled data is scarce.
@@
FixMatch combines pseudo-labeling with consistency regularization through weak and strong data augmentation techniques. For an unlabeled sample , the model first produces a pseudo-label by applying a weak augmentation . This pseudo-label then supervises the corresponding strongly augmented version . The learning objective enforces consistency between the model’s predictions for and . Weak augmentations (e.g., horizontal flips and rotations) preserve the semantic content of the input and facilitate reliable pseudo-labeling. Strong augmentations (e.g., color jitter and cutout) introduce stochastic feature perturbations that can mask or distort key input features [li2024towards]. Despite this, the model is trained to maintain prediction consistency, thereby forcing it to generalize beyond specific feature dependencies.
@@
In the context of malware detection, this mechanism is well suited to learning from features introduced by distribution shifts. Let be a malware sample containing both a known semantic feature and a novel, unseen feature . The weak augmentation retains both features, enabling the model to generate a high-confidence pseudo-label based on . If the strong augmentation masks via stochastic feature perturbation, the detector must rely on to maintain consistency with , thereby incorporating into the learned decision boundary. This implicit adaptation supports drift learning by progressively integrating novel features without explicit annotation. Prior work shows that such training supports better alignment between labeled and unlabeled samples under distribution drift [li2024towards].
@@
In our setting, FixMatch leverages this behavior by combining a supervised loss on labeled samples with an unsupervised consistency loss on confidently pseudo-labeled data. The training dataset at step in the Figure 3 is divided into two sets labeled, at step and unlabeled at step . Let denote the labeled dataset and the unlabeled dataset. The learning objective is:
where is the standard cross-entropy (CE) loss on labeled samples.
@@
Here, denotes the malware detector parameterized by , is the strongly augmented view of the unlabeled input , and is the predicted softmax probability from the weakly augmented version used to obtain pseudo-label . The indicator function filters pseudo-labels based on a confidence threshold . The hyperparameter controls the influence of the unsupervised loss. In our experiments, we set and , following best practices from prior work [sohn2020fixmatch]. This high confidence threshold filters out noisy samples, preventing performance degradation even under significant drift.
4 Overview of CITADEL
@@
In this section, we introduce, CITADEL, a semi-supervised active learning framework for adapting to concept drift in malware detection. The overall workflow is outlined in Algorithm 1 , Algorithm 3, and Algorithm 2, and illustrated in Figure 3.
@@
To support robust learning under evolving feature distributions, CITADEL applies two domain-specific augmentations—Bernoulli bit flip and feature masking— to unlabeled data. As shown in Figure 3, the unlabeled dataset is divided into a weekly perturbed set and a strongly perturbed set . These perturbed variants are incorporated into a unified semi-supervised training objective (in step ) that combines supervised contrastive loss with a standard semi-supervised loss, improving class separation for samples affected by distribution shift. To support adaptation under limited labeling budgets, CITADEL incorporates a multi-criteria sample selection strategy in step that guides the active learning process in step , based on the budget defined in step . This process repeats as new data becomes available, enabling the detector to progressively adapt to concept drift over time.
4.1 Augmentations for Malware Features
@@
CITADEL is built on FixMatch [sohn2020fixmatch], which is originally developed for the computer vision tasks where augmentations such as rotation and color distortion preserve semantic meaning. Malware detection, however, operates over structured feature representations with strict semantic constraints [continual-learning-malware]. In the case of Android malware detection, features are commonly represented as binary vectors derived from static analysis, where each dimension indicates the presence or absence of specific behavioral attributes, permissions, or API calls [arp2014drebin]. These features encode meaningful structural and behavioral properties of malware samples, making image-based augmentations inappropriate in this domain. As a result, effective SSL in this domain requires features perturbations that respect the underlying semantics of malware samples.
To address this challenge, we propose two augmentation strategies tailored for static binary feature vectors—Bernoulli bit flip and Bernoulli feature mask, as shown in Algorithm 1 lines 6–7 and Figure 4. These augmentations are designed to model feature-level variation induced by concept drift. In Android malware, concept drift often manifests as changes in attack strategies rather than objectives. For example, an early variant may abuse SMS permissions to exfiltrate sensitive data via text messages, while a later variant achieves the same objective by encrypted HTTPS requests, activating different static features. Since such drifts do not follow a deterministic pattern [pendlebury2019tesseract], we adopt a randomized perturbation strategy to inject controlled noise into the feature space.
In particular, we use Bernoulli distribution to perturb features since it enables controlled but probabilistic perturbation. Lower probabilities result in mild perturbations that preserve most features, while high probabilities induce stronger variation. This design enables consistency-based regularization within the feature space, encouraging the classifier to rely on features that remain predictive under variation and supporting adaptation to evolving malware behavior.
Prior work such as MORSE [wu2023grim] also applies FixMatch to malware analysis but focuses on handling label noise rather than concept drift. In contrast, our augmentations are explicitly designed to model feature changes arising from evolving attack strategies.
Bernoulli Bit Flip. This augmentation models the feature-level variation across malware variants by probabilistically flipping bits in the binary feature vector. In this representation, a value of 1 indicates the presence of a specific API call or permission, while 0 indicates its absence. Given an input feature vector , we apply independent Bernoulli noise to each feature to control the extent of perturbation. Specifically, we sample a binary mask such that each entry is drawn independently from a Bernoulli distribution with parameter . The perturbed input is obtained via an element-wise XOR operation:
where denotes the element-wise XOR. Small values of introduce subtle perturbations that preserve most features, while larger values of causes substantial perturbations that may degrade semantic coherence.
@@
Bernoulli Feature Mask. In this augmentation, we intend to model scenarios in which observable behaviors are suppressed due to obfuscation or changes in implementation by randomly zeroing out a subset of the features, effectively removing features from the sample. Given an input vector , we sample a binary mask and compute the masked input as:
@@
Weak and Strong Augmentation. Within the CITADEL framework, weak and strong augmentation serve distinct roles. Weak augmentation applies a low perturbation probability preserving class semantics of a sample and is used to generate confident pseudo-labels (shown in Algorithm 1 line: 8). Strong augmentation apples a higher perturbation probability to introduce greater variation during training.
To evaluate the impact of each strategy, we explore the effect of Bernoulli bit flip and Bernoulli feature mask independently on model generalization and concept drift adaptation. For Bernoulli bit flip, we apply weak () and strong () perturbations using flip probabilities and , respectively. For Bernoulli feature mask, we use the same weak and strong masking probabilities, and , respectively. These values are selected based on empirical evaluation (see Appendix C), as they provide the best performance while minimizing label distortion and preventing model performance degradation.
4.2 CITADEL’s Objective Function
@@
The supervised contrastive loss follows the formulation on [khosla2020supervised]. Given a batch of encoded feature vectors and their corresponding labels , the loss is defined as:
@@
4.3 Multi-Criteria Sample Selection
Active learning offers a promising approach to reducing labeling costs in malware detection [chen2023continuous], particularly in the presence of concept drift. Prior work utilizes Mean Absolute Deviation (MAD) to find the distance from its class center using contrastive classifier and applying threshold to find the out-of-distribution (OOD) samples [yang2021cade], Chen-AL [chen2023continuous] uses complex local pseudo loss for selecting uncertain samples; however, both fail to adapt longitudinal and pronounced drift scenario such as in LAMDA [haque2025lamda]. We empirically observe that no single sampling criterion reliably captures all informative instances under such dynamics. To address this, we introduce a multi-criteria sampling strategy (Algorithm 3) based on three complementary indicators: (1) boundary proximity, via the softmax margin; (2) feature displacement,via the -norm in embedding space; and (3) prediction uncertainty, via low-confidence outputs. We hypothesize that their combination enables more robust and stable adaptation in the face of continual malware evolution. Algorithm 2 presents the detailed procedure of active learning method.
Softmax Margin for Boundary Samples. Through softmax prediction margin of the classifier reflects the proximity of a sample to the decision boundary. For an unlabeled input , let denote the classifier’s softmax output vector. The softmax margin is defined as the difference between the top two predicted class probabilities:
where and are the highest and second-highest values in , respectively. A smaller margin indicates greater uncertainty and typically corresponds to samples near the model’s decision boundary. Such samples are more likely to be informative when labeled. Margin-based uncertainty sampling is a well-established strategy in the active learning literature [gal2017deep], and has demonstrated robustness even under distributional shift [ash2019deep].
-Norm for Feature-Space Drift. While the softmax margin captures local uncertainty near the classifier’s decision boundary, it may fail to identify outliers or samples that are semantically distant from the training distribution—a common scenario under concept drift. To address this, we compute the -norm distance between each unlabeled sample and its nearest labeled neighbor in the classifier’s learned feature space.
@@ 4.3 Multi-Criteria Sample Selection
To select an appropriate value, we empirically evaluated the active learning performance for in increments of 0.1. Our results showed only minor performance variation across values. We ultimately select due to its consistent performance across multiple settings and its theoretical grounding in high-dimensional vector spaces [li2002limit, busemann2005geometry].
@@
Samples with low confidence often correspond to ambiguous instances or previously unseen malware classes. Prior work has demonstrated that such samples are strongly correlated with misclassification [hendrycks2017baseline] and can be effectively utilized for active learning under distributional shift [wang2022freematch, zhang2021flexmatch]. In the context of malware detection, this criterion enables the model to prioritize uncertain or suspicious samples that may not be captured by margin-based or distance-based strategies alone.
Sample Selection Strategy. To select the uncertain samples for labeling, we combine three complementary uncertainty criteria: softmax margin (), feature-space distance (), and prediction confidence (). For each unlabeled sample, we first compute all three scores. Since these scores have different ranges and interpretations, we apply min–max normalization to each criterion so that all values lie in the range .
We then align the direction of the scores such that higher values consistently indicate higher uncertainty. Specifically, samples with smaller softmax margins and lower prediction confidence are more uncertain, so we invert their normalized values using and , respectively. After this alignment, we combine the three uncertainty signals by summation to obtain a unified uncertainty score for each unlabeled sample and select the top- samples with the highest scores for labeling. The unified uncertainty score for an unlabeled sample is computed as:
Here, denotes the normalized value. This multi-criteria strategy does not rely on a single uncertainty signal. Instead, it jointly considers boundary ambiguity, feature-space novelty, and low-confidence predictions, which is particularly important under concept drift where predictive confidence alone can be misleading due to shifts in the decision boundary [transcend].
This formulation assigns higher selection scores to samples with low prediction margins, high feature-space distance, and low model confidence—each of which signals potential informativeness under concept drift. The top- samples with the highest scores are selected for labeling.
The sample selection and retraining procedures are detailed in Algorithm 2, lines 10–13, which invoke the semi-supervised training routine in Algorithm 1. Compared to prior active learning approaches in malware detection that rely on a single uncertainty metric [yang2021cade, chen2023continuous], our multi-criteria strategy enables more robust adaptation to drift by jointly targeting decision boundary ambiguity and semantic outliers in the learned feature space.
| Year | APIGraph [api_graph_dataset] | Chen-AZ [chen2023continuous] | LAMDA [haque2025lamda] | MaMaDroid [mamadroid] | ||||
@@
5 Evaluation
In this section, we evaluate CITADEL across multiple settings to assess its effectiveness under concept drift and limited labeling budgets. We first describe the datasets in Section 5.1. then analyzes the behavior of the CITADEL without active learning under varying labeled ratios and feature augmentations in Section 5.2. Section 5.3 compares CITADEL against baseline methods without active learning (AL), while Section 5.4 extends this comparison to the AL framework. Section 5.5 reports the computational complexity.
Additional analyses are provided in the Appendix, including dataset composition statistics (Appendix A), visual analysis comparing CITADEL and Chen-AL performance (Appendix B), hyperparameter sensitivity and augmentation analysis (Appendix C), robustness to label noise (Appendix D), curriculum learning integration (Appendix E), results on additional feature sets such as EMBER (Appendix F), and details of computational resources (Appendix G).